carmelle:

The aquarium has 1 fewer red fish than blue fish. 60% of the fish are blue. How many blue fish are in the aquarium?

Vocaloid:

Let R = # of red fish, B = # blue fish One less red fish than blue —> R = B - 1 60% of the fish are blue —> B/(R+B) = 0.60 Solve the second equation for R in terms of B, plug it into R in the first equation, solve for B

carmelle:

Ok, so what I have so far is: B = 0.6(R+B) - because I multiplied both sides by (R + B) to get rid of the fraction. Now, I'm not sure if I'm supposed to distribute 0.6 into R + B. Is that my next step? If I distribute, won't my final answer be a decimal?

Vocaloid:

After you distribute 0.60 across R and B you can then combine the B terms on one side and the R term on the other side, and finally isolate R by dividing

Vocaloid:

B = (0.60)(R+B) B = 0.60R + 0.60B Subtract 0.60B from both sides 0.40B = 0.60R Solve for R R = (2/3)B Now, since you have this equation as well as the first equation R = B - 1, you can set them equal to each other and solve for B

carmelle:

okok, now i'm understanding, B - 1 = 2/3B + 1 B = 2/3B + 1 1/3B = 1 B = 3 So, the final answer should be, 3 blue fish and 2 red fish. YAY! Thank you so much!!